two different numbers use the same two digits. The sume of the digits is 7, and the difference of the numbers is 27. find the numbers
Answer by htmentor(1343)

two different numbers use the same two digits. The sume of the digits is 7, and the difference of the numbers is 27. find the numbers
=====================
Let t = the tens digit of the 1st number
Let u = the units digit of the 1st number
Then the two numbers can be represented as 10t + u and 10u + t
The difference of the two numbers is 27: 10t + u - (10u + t) = 27
9t + 9u = 27
t - u = 3 [1]
And the sum of the digits = 7: t + u = 7 [2]
Add [1] and [2]:
2t = 10 -> t = 5
So the one's digit = 2
The two numbers are 52 and 25